# Perfect Edge Domination in $P_{6}$-free Graphs and in Graphs Without Efficient Edge Dominating Sets

### v1

@@ 2.1 Efficient and perfect edge domination
Let $G$ be a graph. The vertex and edge sets of $G$ are denoted by $V(G)$ and $E(G)$, respectively. If $uv\in E(G)$, we say that $u$ (resp. $v$) is adjacent to $v$ (resp. $u$); we also say that the edge $uv$ is incident at $u$ (resp. $v$). The set of neighbors of $v$ is denoted by $N_{G}(v)$; i.e., all vertices $u\in V(G)$ adjacent to $v$. When the context is clear, we may write $N(v)$ instead of $N_{G}(v)$. We also define the closed neighborhood of $v$ as $N_{G}[v]=N_{G}(v)\cup\{v\}$. The degree of a vertex $v\in V(G)$ is $d_{G}(v)=|N_{G}(v)|$. A vertex $v$ is called a pendant vertex in $G$ if $d_{G}(v)=1$. Two edges $e$ and $e^{\prime}$ are adjacent if they are incident at the same vertex. If $S\subseteq V(G)$, the subgraph induced by $S$ is denote by $G[S]$, and GSG-S stands for $G[V(G)\setminus S]$. Moreover, for a vertex $v\in V(G)$, we define $N_{G}^{S}(v)=N_{G}(v)\cap S$. If $H$ is a subgraph of $G$, we let $N_{H}(v)=N^{V(H)}_{G}(v)$. A subgraph $H$ of $G$ is called a spanning subgraph of $G$ if $V(H)=V(G)$, and it is called a spanning tree if $H$ is also a tree (i.e., a connected and acyclic graph).
@@
An edge $uv$ dominates itself and any other edge adjacent to it. A subset $P\subseteq E(G)$ is a perfect edge dominating set (PED-set) if every edge of $E(G)\setminus P$ is dominated by exactly one edge of $P$.
@@
An assignment of one of the three possible colors to each vertex of $G$ will be called a coloring of $G$, and we refer to $(B,Y,W)$ as the $3$-coloring associated with $P$. It is easy to see that the following conditions hold:
@@
Besides, a $3$-coloring $(B,Y,W)$ of a graph $G$ satisfying the four above conditions is associated with a perfect edge dominating set of $G$, consisting precisely of those edges having non-white endpoints (i.e., vertices with a non-white color assigned). A coloring is called partial if only some vertices of $G$ have assigned colors; otherwise, it is called total. A partial coloring is valid if no two white vertices are adjacent, no yellow vertex has more than one non-white neighbor, and no black vertex has white neighbors. A total coloring is valid if no two white vertices are adjacent, every yellow vertex has exactly one non-white neighbor, and no black vertex $v$ has white neighbors, and $d_{G}(v)\geq 2$.
Notice that, if $H$ is a subgraph of a graph $G$, and $\pi$ is a valid $3$-coloring of $G$; i.e., a function assigning to each vertex in $V(G)$ a color black (b), yellow (y), and white (w), then $\pi$ restricted to $V(H)$ is a valid partial $3$-coloring of $H$. It is said that a valid partial $3$-coloring $\pi$ of a subgraph $H$ of $G$ can be extended to a valid $3$-coloring of $G$, if $G$ has a valid $3$-coloring $\pi^{\prime}$ such that, restricted to $V(H)$, it agrees with $\pi$, meaning that $\pi^{\prime}(u)=\pi(u)$, for all $u\in V(H)\subseteq V(G)$. If $\pi$ generates a PED-set of minimum cardinality, we say that $\pi$ is an optimum valid $3$-coloring.
@@
Let $(B,Y,W)$ be the valid $3$-coloring of the weighted graph $G$ induced by an edge dominating set $P$. Define the function $\psi$ as $\psi(u)=\displaystyle\sum_{v\in N_{G}(u)}\omega(uv)$ for every $u\in V(G)$. Then, the weight of the PED-set $P$ in $G$ satisfies
@@ 2.2 $P_{6}$-free graphs
Our cubic-time algorithm in Section 4 for finding a PED-set of minimum cardinality for $P_{6}$-free graphs relies on considering all possible valid partial $3$-colorings for either the dominating induced $C_{6}$ or the dominating complete bipartite subgraph.
@@ 3 DIM-less graphs
If $G$ has a DIM, then $v$ is either white or yellow. If $v$ is white (resp. yellow), we can complete the $3$-coloring to $G^{\prime}$ by using the coloring of Fig.1, changing $v_{1}$ from gray to yellow (resp. white), and $v_{2}$ from gray to white (resp. yellow). Clearly, the resulting coloring corresponds to a non-trivial PED-set of $G^{\prime}$.
@@ 4.1 Graphs with a dominating induced $C_{6}$
Let us begin by considering connected $P_{6}$-free graphs that have a dominating induced $C_{6}$. The following lemma outlines the possible valid partial $3$-colorings for such a cycle.
@@
Let $G$ be a graph. All valid partial $3$-colorings $(B,Y,W)$ for an induced $C_{6}$ in $G$ are depicted in Fig. 2.
@@
Let $G$ be a graph with a valid $3$-coloring $(B,Y,W)$, and let $C=(v_{1},v_{2},v_{3},v_{4},v_{5},v_{6})$ be an induced cycle in $G$. Since a cycle has no pendant vertices, if $v_{i}\notin W$ for all $1\leq i\leq 6$, then $v_{i}\in B$ for each $i$. This corresponds to the valid partial $3$-coloring shown in case (a).
@@
  • Case 3: If $v_{3}\in B$, then $v_{4}\in B$ or $v_{4}\in Y$, giving place to the valid partial $3$-coloring (b) or (c) of $C$, respectively.
Hence, all valid partial $3$-colorings of the induced $C_{6}$ are as depicted. ∎
@@
Let $G$ be a connected $P_{6}$-free graph, and let $C$ be a dominating induced $C_{6}$ in $G$. Then $|N_{C}(v)|\geq 2$ for every $v\in V(G)\setminus V(C)$. Moreover, if $(B,Y,W)$ is a valid $3$-coloring of $G$, then one of the following conditions holds:
  • The corresponding partial $3$-coloring of $C$ is of type (a), with $B=V(G)$ and $Y=W=\emptyset$. This valid 3-coloring corresponds to the trivial PED-set.
  • The corresponding partial $3$-coloring of $C$ is of type (b), $c$ is the unique black vertex of $C$ and $d\in V(C)$ is the unique yellow vertex with only white neighbors in $V(C)$, then:
@@
  • The corresponding partial $3$-coloring of $C$ is of type (c) with:
@@
  • The corresponding partial $3$-coloring of $C$ is of type (d) with:
@@
  • The corresponding partial $3$-coloring of $C$ is of type (e) with:
@@
Now, consider each valid partial $3$-coloring type of $C$:
@@
Figure 2: Five possible valid partial $3$-colorings of an induced dominating $C_{6}$
@@ 4.2 Graphs with a dominating complete bipartite subgraph
Now, let us examine a connected $P_{6}$-free graph $G$ that contains a dominating, though not necessarily induced, complete bipartite subgraph $K$, where $K$ includes $K_{p,q}$ as a spanning subgraph with a bipartition $\{R,S\}$, where $|R|=p$ and $|S|=q$. We will explore all possible valid partial $3$-colorings for $K$. It is no hard to see that any dominating complete bipartite subgraph can be extended to a maximal one in linear-time. Hence, we can assume that $K$ is maximal, assuming this additional cost.
@@ 4.2.1 $V(K)\cap B=\emptyset$
First, we will consider the case in which $K$ has a valid partial $3$-coloring with no black vertices. Let us start by considering the case in which $G$ is also a bipartite graph.
@@
Let $G$ be a bipartite connected $P_{6}$-free graph with at least two vertices, and let $\{M,N\}$ be its bipartition. If $G$ admits a valid $3$-coloring $(B,Y,W)$ such that $N\subseteq Y$, then there exists a vertex in $M$ that is adjacent to every vertex in $N$.
@@
Assume now that $p\geq 2$. Suppose that $(B,Y,W)$ is a valid $3$-coloring of $G$, and let $u\in U$ be a non-white vertex. Then, there can be no other non-white vertex in $M$; otherwise, some vertex in $V$ would be adjacent to more than one non-white neighbor, contradicting the properties of yellow vertices. Therefore, $u$ must be adjacent to every vertex in $N$.
@@
Suppose, towards a contradiction, that no vertex in $M$ is adjacent to all vertices in $N$. Hence, there exists a vertex $d_{1}\in N\setminus V$. Since $d_{1}\in Y$, it can be adjacent to at most one of $c_{1}$ and $c_{2}$. Suppose, without loss of generality, that $d_{1}$ is adjacent only to $c_{1}$. Since $p+q$ is maximum, there exists $u\in U$ such that $u$ is non-adjacent to $d_{1}$. Consequently, the set $\{d_{1},c_{1},v_{1},u,v_{2},c_{2}\}$ induces $P_{6}$, contradicting the assumption that $G$ is $P_{6}$-free.
@@
In all cases, assuming that no vertex in $M$ is adjacent to every vertex in $N$ leads to a contradiction. Therefore, such a vertex must exist. ∎
The following lemma describes a possible valid partial $3$-coloring of $K$ when $V(K)\cap B=\emptyset$. Recall that $\{R,\,S\}$ denotes the bipartition of $K$.
@@
Let $G$ be a $P_{6}$-free graph that contains a dominating (not necessarily induced) complete bipartite subgraph $K$. If $G$ admits a valid $3$-coloring $(B,Y,W)$ such that $V(K)\cap B=\emptyset$, then exactly one of the following conditions holds:
@@
The following lemma describes how all possible valid partial $3$-coloring for $K$ can be extended to a valid $3$-coloring for $G$ when $|R\cap Y|=|S\cap Y|=1$.
Lemma 4.4.
Let $G$ be a $P_{6}$-free graph containing a maximal dominating (not necessarily induced) complete bipartite subgraph $K$. If $G$ has a valid $3$-coloring $(B,Y,W)$ such that $|R\cap Y|=|S\cap Y|=1$, then:
@@
        Let $G^{\prime}=G-(R\cup Y^{\prime}\cup W^{\prime})$. Any pair of vertices $x,x^{\prime}\in S\setminus W^{\prime}$ verifies $N_{G^{\prime}}(x)\cap N_{G^{\prime}}(x^{\prime})=\emptyset$. Let $W^{\prime\prime}=\{s\in S\setminus W^{\prime}:N_{G^{\prime}}(s)\mbox{ contains two adjacent vertices}\}$, and let $Y^{\prime\prime}=\{x\in V(G)\setminus(V(K)\cup Y^{\prime}):N^{W^{\prime\prime}}(x)\neq\emptyset\}$. Then $S^{\prime}=S\setminus(W^{\prime}\cup W^{\prime\prime})\neq\emptyset$ for some vertex $s\in S^{\prime}\cap Y$, $V(G)\setminus(V(K)\cup Y^{\prime}\cup Y^{\prime\prime}\cup N(s))\cup\{s\}\subseteq Y$, and $(V(G)\setminus(V(K)\cup Y^{\prime}\cup Y^{\prime\prime})\cap N(s))\cup(S^{\prime}\setminus\{s\})\subseteq W$. In addition, $V(G)\setminus(V(K)\cup N(s))$ induces a $1$-regular subgraph in $G$.
@@
Let $x\in W^{\prime\prime}$ with two adjacent neighbors $y$ and $z$ in $V(G)\setminus(V(K)\cup Y^{\prime})$. Suppose that $x$ is a yellow vertex. Thus, $y$ must be yellow and $z$ must be white, or vice versa. Additionally, $x$ is a yellow vertex with two non-white neighbors, $y$ (or $z$) and $r$, contradicting that the $3$-coloring is admissible. Therefore, $W^{\prime\prime}\subseteq W$ and $Y^{\prime\prime}\subseteq Y$. The rest of this part of the statement follows by the definition of admissible $3$-coloring.
Suppose that $E(G[S])=\{ss^{\prime}\}$. Since $s$ and $s^{\prime}$ are adjacent to the yellow vertex $r$, one of them must be a yellow vertex. The color of the vertices in $V(G)\setminus V(K)$ depends on they are adjacent or non-adjacent to this yellow vertices. Therefore, there are at most two valid total $3$-colorings.
Suppose now that $|E(G[S])|\geq 2$. Since $r$ is a yellow vertex, $G[S]$ is isomorphic to $K_{1,g}+hK_{1}$ with $g+h=|S|$. Additionally, the vertex $x$ of maximum degree in $G[S]$ must be the yellow vertex of $S$. Consequently, there is only one valid $3$-coloring.
Finally, it is clear that $B=\emptyset$. Therefore, each valid total $3$-coloring is a DIM. ∎
Given a connected $P_{6}$-free graph $G$ with a dominating (not necessarily induced) complete bipartite subgraph $K$, and a valid $3$-coloring $(B,Y,W)$ such that $R\subseteq W$ and $S\subseteq Y$, we use $X$ to denote the subset of vertices in $V(G)\setminus V(K)$ with at least one neighbor in $R$. We denote the subgraph $G-(V(K)\cup X)$ by $G_{0}$. Note that $S$ dominates $V(G_{0})$. Let $C$ be a connected component of $G_{0}$. When $C$ is bipartite, we denote the bipartition of $C$ by $\{U_{C},V_{C}\}$.
@@
Let $G$ be a connected $P_{6}$-free graph that contains a dominating (not necessarily induced) complete bipartite subgraph $K$. If $G$ admits a valid $3$-coloring $(B,Y,W)$ such that $R\subseteq W$ and $S\subseteq Y$, then the following conditions hold:
@@
Let $\pi$ be a valid $3$-coloring assignment of $G$.
Statement $1.$ follows from the definition of a valid $3$-coloring and the fact that $R\subseteq W$.
@@
By $S_{1}$ (resp. $X_{1})$, we refer to the subset of vertices $v\in S$ (resp. $v\in X)$ such that $d_{G[X\cup S]}(v)=1$. Let $S_{2}=S\setminus S_{1}$ (resp. $X_{2}=X\setminus X_{1}$). Note that $X_{2}\cup S_{2}$ is an independent set, since $d_{G[X\cup S]}(v)=0$ for all $v\in X_{2}\cup S_{2}$. Let $C$ be a non-trivial bipartite connected component of $G_{0}$ that has at least one vertex $x$ with a neighbor in $S_{2}$. We say that $C$ is of type I if there exists a vertex $v\in S_{2}$ such that $v$ has a neighbor $c$ and a non-neighbor $d$, both in $V(C)$, and $cd$ is an edge of $E(C)$. Such vertex $v$ is called an indicator of $C$. The component $C$ is called of type II if $|V(C)|\geq 2$ and every vertex in $S_{2}$ with a neighbor in this component is adjacent to each vertex of $V(C)$.
@@
Let $G$ be a connected $P_{6}$-free graph having a dominating (not necessarily induced) complete bipartite subgraph $K$ with $R\subseteq W$ and $S\subseteq Y$. If $N_{G}(v)\cap V(G_{0})=\emptyset$ for some $v\in X_{2}\cup S_{2}$, then $G$ has no valid $3$-coloring.
@@
Suppose, towards a contradiction, that $G$ has a valid $3$-coloring, and $v$ is a vertex in $X_{2}\cup S_{2}$ such that $N_{G}(v)\cap V(G_{0})=\emptyset$. If $u\in N_{G}(v)$, then $u\not\in V(G_{0})$; i.e., $u\in V(K)\cup X=R\cup S\cup X$. Furthermore, $u\not\in X\cup S$ because, according to the definition of $X_{2}$ and $S_{2}$, it follows that $d_{G[X\cup S]}(v)=0$. Hence, $u\in R$. Since $R\subseteq W$, $u\in W$. Thus, as $X_{2}\cup S_{2}\subseteq Y$, $v$ is a yellow vertex with all its neighbors being white vertices. Therefore, $G$ has no valid 3-coloring, contradicting the assumption that $G$ has a valid $3$-coloring. ∎
The following lemma ensures that the colors assigned to the vertices in a connected component of type I within a valid $3$-coloring for $G$ also determine the colors of each vertex in any other component of type I.
@@
The below lemma characterizes the component of type II of $G_{0}$ when $G$ has a valid $3$-coloring.
@@
Let $G$ be a connected $P_{6}$-free graph with a dominating (not necessarily induced) complete bipartite subgraph $K$ where $R\subseteq W$ and $S\subseteq Y$. If $G$ has a valid $3$-coloring, and $C$ is a connected component of $G_{0}$ of type II, then either $|U_{C}|=1$, or $|V_{C}|=1$.

### v2

@@ 2.1 Efficient and perfect edge domination
Let $G$ be a graph. The vertex and edge sets of $G$ are denoted by $V(G)$ and $E(G)$, respectively. If $uv\in E(G)$, we say that $u$ (resp. $v$) is adjacent to $v$ (resp. $u$); we also say that the edge uvuv is incident at $u$ (resp. $v$). The set of neighbors of $v$ is denoted by $N_{G}(v)$; i.e., all vertices $u\in V(G)$ adjacent to $v$. When the context is clear, we may write $N(v)$ instead of $N_{G}(v)$. We also define the closed neighborhood of $v$ as $N_{G}[v]=N_{G}(v)\cup\{v\}$. The degree of a vertex $v\in V(G)$ is $d_{G}(v)=|N_{G}(v)|$. A vertex $v$ is called a pendant vertex in $G$ if $d_{G}(v)=1$. Two edges $e$ and $e^{\prime}$ are adjacent if they are incident at the same vertex. If $S\subseteq V(G)$, the subgraph induced by $S$ is denote by $G[S]$, and GSG-S stands for $G[V(G)\setminus S]$. Moreover, for a vertex $v\in V(G)$, we define $N_{G}^{S}(v)=N_{G}(v)\cap S$. If $H$ is a subgraph of $G$, we let $N_{H}(v)=N^{V(H)}_{G}(v)$. A subgraph $H$ of $G$ is called a spanning subgraph of $G$ if $V(H)=V(G)$, and it is called a spanning tree if $H$ is also a tree (i.e., a connected and acyclic graph).
@@
An edge uvuv dominates itself and any other edge adjacent to it. A subset $P\subseteq E(G)$ is a perfect edge dominating set (PED-set) if every edge of $E(G)\setminus P$ is dominated by exactly one edge of $P$.
@@
An assignment of one of the three possible colors to each vertex of $G$ will be called a coloring of $G$, and we refer to $(B,Y,W)$ as the 33-coloring associated with $P$. It is easy to see that the following conditions hold:
@@
Besides, a 33-coloring $(B,Y,W)$ of a graph $G$ satisfying the four above conditions is associated with a perfect edge dominating set of $G$, consisting precisely of those edges having non-white endpoints (i.e., vertices with a non-white color assigned). A coloring is called partial if only some vertices of $G$ have assigned colors; otherwise, it is called total. A partial coloring is valid if no two white vertices are adjacent, no yellow vertex has more than one non-white neighbor, and no black vertex has white neighbors. A total coloring is valid if no two white vertices are adjacent, every yellow vertex has exactly one non-white neighbor, and no black vertex $v$ has white neighbors, and $d_{G}(v)\geq 2$.
Notice that, if $H$ is a subgraph of a graph $G$, and $\pi$ is a valid 33-coloring of $G$; i.e., a function assigning to each vertex in $V(G)$ a color black (b), yellow (y), and white (w), then $\pi$ restricted to $V(H)$ is a valid partial 33-coloring of $H$. It is said that a valid partial 33-coloring $\pi$ of a subgraph $H$ of $G$ can be extended to a valid 33-coloring of $G$, if $G$ has a valid 33-coloring $\pi^{\prime}$ such that, restricted to $V(H)$, it agrees with $\pi$, meaning that $\pi^{\prime}(u)=\pi(u)$, for all $u\in V(H)\subseteq V(G)$. If $\pi$ generates a PED-set of minimum cardinality, we say that $\pi$ is an optimum valid 33-coloring.
@@
Let $(B,Y,W)$ be the valid 33-coloring of the weighted graph $G$ induced by an edge dominating set $P$. Define the function $\psi$ as $\psi(u)=\displaystyle\sum_{v\in N_{G}(u)}\omega(uv)$ for every $u\in V(G)$. Then, the weight of the PED-set $P$ in $G$ satisfies
@@ 2.2 $P_{6}$-free graphs
Our cubic-time algorithm in Section 4 for finding a PED-set of minimum cardinality for $P_{6}$-free graphs relies on considering all possible valid partial 33-colorings for either the dominating induced $C_{6}$ or the dominating complete bipartite subgraph.
@@ 3 DIM-less graphs
If $G$ has a DIM, then $v$ is either white or yellow. If $v$ is white (resp. yellow), we can complete the 33-coloring to $G^{\prime}$ by using the coloring of Fig.1, changing $v_{1}$ from gray to yellow (resp. white), and $v_{2}$ from gray to white (resp. yellow). Clearly, the resulting coloring corresponds to a non-trivial PED-set of $G^{\prime}$.
@@ 4.1 Graphs with a dominating induced $C_{6}$
Let us begin by considering connected $P_{6}$-free graphs that have a dominating induced $C_{6}$. The following lemma outlines the possible valid partial 33-colorings for such a cycle.
@@
Let $G$ be a graph. All valid partial 33-colorings $(B,Y,W)$ for an induced $C_{6}$ in $G$ are depicted in Fig. 2.
@@
Let $G$ be a graph with a valid 33-coloring $(B,Y,W)$, and let $C=(v_{1},v_{2},v_{3},v_{4},v_{5},v_{6})$ be an induced cycle in $G$. Since a cycle has no pendant vertices, if $v_{i}\notin W$ for all $1\leq i\leq 6$, then $v_{i}\in B$ for each ii. This corresponds to the valid partial 33-coloring shown in case (a).
@@
  • Case 3: If $v_{3}\in B$, then $v_{4}\in B$ or $v_{4}\in Y$, giving place to the valid partial 33-coloring (b) or (c) of $C$, respectively.
Hence, all valid partial 33-colorings of the induced $C_{6}$ are as depicted. ∎
@@
Let $G$ be a connected $P_{6}$-free graph, and let $C$ be a dominating induced $C_{6}$ in $G$. Then $|N_{C}(v)|\geq 2$ for every $v\in V(G)\setminus V(C)$. Moreover, if $(B,Y,W)$ is a valid 33-coloring of $G$, then one of the following conditions holds:
  • The corresponding partial 33-coloring of $C$ is of type (a), with $B=V(G)$ and $Y=W=\emptyset$. This valid 3-coloring corresponds to the trivial PED-set.
  • The corresponding partial 33-coloring of $C$ is of type (b), $c$ is the unique black vertex of $C$ and $d\in V(C)$ is the unique yellow vertex with only white neighbors in $V(C)$, then:
@@
  • The corresponding partial 33-coloring of $C$ is of type (c) with:
@@
  • The corresponding partial 33-coloring of $C$ is of type (d) with:
@@
  • The corresponding partial 33-coloring of $C$ is of type (e) with:
@@
Now, consider each valid partial 33-coloring type of $C$:
@@
Figure 2: Five possible valid partial 33-colorings of an induced dominating $C_{6}$
@@ 4.2 Graphs with a dominating complete bipartite subgraph
Now, let us examine a connected $P_{6}$-free graph $G$ that contains a dominating, though not necessarily induced, complete bipartite subgraph $K$, where $K$ includes $K_{p,q}$ as a spanning subgraph with a bipartition $\{R,S\}$, where $|R|=p$ and $|S|=q$. We will explore all possible valid partial 33-colorings for $K$. It is no hard to see that any dominating complete bipartite subgraph can be extended to a maximal one in linear-time. Hence, we can assume that $K$ is maximal, assuming this additional cost.
@@ 4.2.1 $V(K)\cap B=\emptyset$
First, we will consider the case in which $K$ has a valid partial 33-coloring with no black vertices. Let us start by considering the case in which $G$ is also a bipartite graph.
@@
Let $G$ be a bipartite connected $P_{6}$-free graph with at least two vertices, and let $\{M,N\}$ be its bipartition. If $G$ admits a valid 33-coloring $(B,Y,W)$ such that $N\subseteq Y$, then there exists a vertex in $M$ that is adjacent to every vertex in NN.
@@
Assume now that $p\geq 2$. Suppose that $(B,Y,W)$ is a valid 33-coloring of $G$, and let $u\in U$ be a non-white vertex. Then, there can be no other non-white vertex in $M$; otherwise, some vertex in $V$ would be adjacent to more than one non-white neighbor, contradicting the properties of yellow vertices. Therefore, $u$ must be adjacent to every vertex in NN.
@@
Suppose, towards a contradiction, that no vertex in $M$ is adjacent to all vertices in NN. Hence, there exists a vertex $d_{1}\in N\setminus V$. Since $d_{1}\in Y$, it can be adjacent to at most one of $c_{1}$ and $c_{2}$. Suppose, without loss of generality, that $d_{1}$ is adjacent only to $c_{1}$. Since $p+q$ is maximum, there exists $u\in U$ such that $u$ is non-adjacent to $d_{1}$. Consequently, the set $\{d_{1},c_{1},v_{1},u,v_{2},c_{2}\}$ induces $P_{6}$, contradicting the assumption that $G$ is $P_{6}$-free.
@@
In all cases, assuming that no vertex in $M$ is adjacent to every vertex in NN leads to a contradiction. Therefore, such a vertex must exist. ∎
The following lemma describes a possible valid partial 33-coloring of $K$ when $V(K)\cap B=\emptyset$. Recall that $\{R,\,S\}$ denotes the bipartition of $K$.
@@
Let $G$ be a $P_{6}$-free graph that contains a dominating (not necessarily induced) complete bipartite subgraph $K$. If $G$ admits a valid 33-coloring $(B,Y,W)$ such that $V(K)\cap B=\emptyset$, then exactly one of the following conditions holds:
@@
The following lemma describes how all possible valid partial 33-coloring for $K$ can be extended to a valid 33-coloring for $G$ when $|R\cap Y|=|S\cap Y|=1$.
Lemma $4.$
Let $G$ be a $P_{6}$-free graph containing a maximal dominating (not necessarily induced) complete bipartite subgraph $K$. If $G$ has a valid 33-coloring $(B,Y,W)$ such that $|R\cap Y|=|S\cap Y|=1$, then:
@@
        Let $G^{\prime}=G-(R\cup Y^{\prime}\cup W^{\prime})$. Any pair of vertices $x,x^{\prime}\in S\setminus W^{\prime}$ verifies $N_{G^{\prime}}(x)\cap N_{G^{\prime}}(x^{\prime})=\emptyset$. Let $W^{\prime\prime}=\{s\in S\setminus W^{\prime}:N_{G^{\prime}}(s)\mbox{ contains two adjacent vertices}\}$, and let $Y^{\prime\prime}=\{x\in V(G)\setminus(V(K)\cup Y^{\prime}):N^{W^{\prime\prime}}(x)\neq\emptyset\}$. Then $S^{\prime}=S\setminus(W^{\prime}\cup W^{\prime\prime})\neq\emptyset$ for some vertex $s\in S^{\prime}\cap Y$, $V(G)\setminus(V(K)\cup Y^{\prime}\cup Y^{\prime\prime}\cup N(s))\cup\{s\}\subseteq Y$, and $(V(G)\setminus(V(K)\cup Y^{\prime}\cup Y^{\prime\prime})\cap N(s))\cup(S^{\prime}\setminus\{s\})\subseteq W$. In addition, $V(G)\setminus(V(K)\cup N(s))$ induces a 11-regular subgraph in $G$.
@@
Let $x\in W^{\prime\prime}$ with two adjacent neighbors $y$ and $z$ in $V(G)\setminus(V(K)\cup Y^{\prime})$. Suppose that $x$ is a yellow vertex. Thus, $y$ must be yellow and $z$ must be white, or vice versa. Additionally, $x$ is a yellow vertex with two non-white neighbors, $y$ (or $z$) and $r$, contradicting that the 33-coloring is admissible. Therefore, $W^{\prime\prime}\subseteq W$ and $Y^{\prime\prime}\subseteq Y$. The rest of this part of the statement follows by the definition of admissible 33-coloring.
Suppose that $E(G[S])=\{ss^{\prime}\}$. Since $s$ and $s^{\prime}$ are adjacent to the yellow vertex $r$, one of them must be a yellow vertex. The color of the vertices in $V(G)\setminus V(K)$ depends on they are adjacent or non-adjacent to this yellow vertices. Therefore, there are at most two valid total 33-colorings.
Suppose now that $|E(G[S])|\geq 2$. Since $r$ is a yellow vertex, $G[S]$ is isomorphic to $K_{1,g}+hK_{1}$ with $g+h=|S|$. Additionally, the vertex $x$ of maximum degree in $G[S]$ must be the yellow vertex of $S$. Consequently, there is only one valid 33-coloring.
Finally, it is clear that $B=\emptyset$. Therefore, each valid total 33-coloring is a DIM. ∎
Given a connected $P_{6}$-free graph $G$ with a dominating (not necessarily induced) complete bipartite subgraph $K$, and a valid 33-coloring $(B,Y,W)$ such that $R\subseteq W$ and $S\subseteq Y$, we use $X$ to denote the subset of vertices in $V(G)\setminus V(K)$ with at least one neighbor in $R$. We denote the subgraph $G-(V(K)\cup X)$ by $G_{0}$. Note that $S$ dominates $V(G_{0})$. Let $C$ be a connected component of $G_{0}$. When $C$ is bipartite, we denote the bipartition of $C$ by $\{U_{C},V_{C}\}$.
@@
Let $G$ be a connected $P_{6}$-free graph that contains a dominating (not necessarily induced) complete bipartite subgraph $K$. If $G$ admits a valid 33-coloring $(B,Y,W)$ such that $R\subseteq W$ and $S\subseteq Y$, then the following conditions hold:
@@
Let $\pi$ be a valid 33-coloring assignment of $G$.
Statement $1.$ follows from the definition of a valid 33-coloring and the fact that $R\subseteq W$.
@@
By $S_{1}$ (resp. $X_{1})$, we refer to the subset of vertices $v\in S$ (resp. $v\in X)$ such that $d_{G[X\cup S]}(v)=1$. Let $S_{2}=S\setminus S_{1}$ (resp. $X_{2}=X\setminus X_{1}$). Note that $X_{2}\cup S_{2}$ is an independent set, since $d_{G[X\cup S]}(v)=0$ for all $v\in X_{2}\cup S_{2}$. Let $C$ be a non-trivial bipartite connected component of $G_{0}$ that has at least one vertex $x$ with a neighbor in $S_{2}$. We say that $C$ is of type I if there exists a vertex $v\in S_{2}$ such that $v$ has a neighbor $c$ and a non-neighbor $d$, both in $V(C)$, and cdcd is an edge of $E(C)$. Such vertex $v$ is called an indicator of $C$. The component $C$ is called of type II if $|V(C)|\geq 2$ and every vertex in $S_{2}$ with a neighbor in this component is adjacent to each vertex of $V(C)$.
@@
Let $G$ be a connected $P_{6}$-free graph having a dominating (not necessarily induced) complete bipartite subgraph $K$ with $R\subseteq W$ and $S\subseteq Y$. If $N_{G}(v)\cap V(G_{0})=\emptyset$ for some $v\in X_{2}\cup S_{2}$, then $G$ has no valid 33-coloring.
@@
Suppose, towards a contradiction, that $G$ has a valid 33-coloring, and $v$ is a vertex in $X_{2}\cup S_{2}$ such that $N_{G}(v)\cap V(G_{0})=\emptyset$. If $u\in N_{G}(v)$, then $u\not\in V(G_{0})$; i.e., $u\in V(K)\cup X=R\cup S\cup X$. Furthermore, $u\not\in X\cup S$ because, according to the definition of $X_{2}$ and $S_{2}$, it follows that $d_{G[X\cup S]}(v)=0$. Hence, $u\in R$. Since $R\subseteq W$, $u\in W$. Thus, as $X_{2}\cup S_{2}\subseteq Y$, $v$ is a yellow vertex with all its neighbors being white vertices. Therefore, $G$ has no valid 3-coloring, contradicting the assumption that $G$ has a valid 33-coloring. ∎
The following lemma ensures that the colors assigned to the vertices in a connected component of type I within a valid 33-coloring for $G$ also determine the colors of each vertex in any other component of type I.
@@
The below lemma characterizes the component of type II of $G_{0}$ when $G$ has a valid 33-coloring.
@@
Let $G$ be a connected $P_{6}$-free graph with a dominating (not necessarily induced) complete bipartite subgraph $K$ where $R\subseteq W$ and $S\subseteq Y$. If $G$ has a valid 33-coloring, and $C$ is a connected component of $G_{0}$ of type II, then either $|U_{C}|=1$, or $|V_{C}|=1$.